OpenStudy (anonymous):

use differentials or linear approximation to find an approximated value for 1/∛(8.2)

OpenStudy (anonymous):

@JoelSjogren

myininaya (myininaya):

the cube root of a 8 is easy to find right? so... let \[f(x)=\frac{1}{\sqrt[3]{x}}\] Then find f'(x) Then find f'(8) Then say f(x) is approximately f'(8)(x-8)+f(8) let me know when you get this far

OpenStudy (anonymous):

A hint: before differentiating \(f\), rewrite it as \(f(x) = \cfrac{1}{x^{1/3}} = x^{-1/3}\).

OpenStudy (anonymous):

Now that I think of it, maybe that's unnecessary. What do you think?

myininaya (myininaya):

I think it is a groovy hint.

myininaya (myininaya):

Now you can apply power rule in that form. I prefer that over quotient rule.

myininaya (myininaya):

Basically we are finding the tangent line to that curve at x=8 For values near 8 we should get some pretty close approximations but as we move further from the point that we found the tangent line that particular line actually becomes a bad approximation to use

OpenStudy (anonymous):

f'(x)= -1/3x^4/3

myininaya (myininaya):

that x^4/3 is also in the denominator right?

myininaya (myininaya):

-1/(3x^(4/3))

OpenStudy (anonymous):

yes and f'(8)= -1/48

myininaya (myininaya):

great so far

OpenStudy (anonymous):

f(x) is approximately (-1/48)(x-8)+(1/2)

myininaya (myininaya):

and what do we want to find we want to find 1/cuberoot(8.2) but f(x)=1/cuberoot(x) so if 1/cuberoot(x) is approx (-1/48)(x-8)+1/2 then what do we want to plug in for x?

OpenStudy (anonymous):

f(x)=(-1/48)x+(1/12)

OpenStudy (anonymous):

you'd plug in 8.2?

myininaya (myininaya):

right and in the function approximation you wrote i think that 1/12 should be a 1/2 :)

myininaya (myininaya):

oh you distribute sorry

OpenStudy (anonymous):

so f(x)=(-1/48)(8.2)+(1/12) correct?

myininaya (myininaya):

is it 2/3 instead of 1/12 though?

OpenStudy (anonymous):

youre right. i multiplied 1/6 and 1/2 instead of adding them

myininaya (myininaya):

and f(x) isn't approx that but f(8.2) is approx that so f(8.2) approx= -1/48*(8.2)+2/3

myininaya (myininaya):

and then the rest is algebra

myininaya (myininaya):

the cool thing about this way is you can find 1/cuberoot(8.2) without using a calculator

myininaya (myininaya):

or at least approximate it

OpenStudy (anonymous):

okay so i got 119/240 but how would you find 1/cuberoot(8.2) without a calculator?

myininaya (myininaya):

i meant approximately by using the above way

OpenStudy (aum):

I think the 2/3 at the end should be a 1/2.

myininaya (myininaya):

119/240 is approximately 1/cuberoot(8.2)

myininaya (myininaya):

no she distributed @aum

myininaya (myininaya):

she did have \[f(x) \approx \frac{-1}{48}(x-8)+\frac{1}{2}\]

myininaya (myininaya):

\[f(x) \approx \frac{-1}{48}x +\frac{8}{48}+\frac{1}{2}\]

OpenStudy (anonymous):

you have to distribute, dont you?

OpenStudy (aum):

\(\Large f'(a) = \lim_{x \rightarrow a}\frac{f(x)-f(a)}{x-a} \) When x is very close to 'a', we can ignore the limit and use the following approximation: \(\Large f'(a) \approx \frac{f(x)-f(a)}{x-a} \) \(\Large f(x) \approx f'(a)*(x-a) + f(a) \) \(\Large f(8.2) \approx f'(8)*(8.2-8.0) + f(8) \)

myininaya (myininaya):

\[f(x) \approx \frac{-1}{48}x+\frac{1}{6}+\frac{1}{2}=\frac{-1}{48}x+\frac{2}{12}+\frac{6}{12}=\frac{-1}{48}x+\frac{8}{12} \\ \] Therefore \[f(x) \approx \frac{-1}{48}x+\frac{2}{3}\]

myininaya (myininaya):

tiff you did everything right you found an approximation to the value you were asked to approximate using the linearzation method

OpenStudy (anonymous):

okay, thank you so much for the help! math definitely is NOT my strong suite

myininaya (myininaya):

you can see your approximation is pretty close since 119/240 is .4958333333333333333333(repeating) and the actual value using the calculator which is just approximation anyways is .495901 these two values are pretty close

myininaya (myininaya):

you did pretty good you actually did the power rule which a lot people struggle with I don't think you are as bad as you think